40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 11 11:33:24 2011, josephb, Update, CDS, Fixed mistake in Matrix of Filter banks naming convention 
    Reply  Sun Mar 13 04:00:53 2011, rana, Update, CDS, ETMY Sim work 
       Reply  Sun Mar 13 16:46:48 2011, josephb, Update, CDS, ETMY Sim work 
Message ID: 4409     Entry time: Sun Mar 13 16:46:48 2011     In reply to: 4408
Author: josephb 
Type: Update 
Category: CDS 
Subject: ETMY Sim work 

4. The blue Output Filters  section has been changed to agree with the new filter of matrices row, column labeling.  My fault for not testing it and realizing it was broken.  The change was made in /opt/rtcds/caltech/c1/medm/master/C1SUS_DEFAULTNAME.adl and then ,/generate_master_screens.py was run, updating all the screens.

5.  I have swapped the logic for the sensor filter banks (ULSEN, URSEN, etc).  It now sends a "1" to the Binary Output board controlling the OSEM analog whitening when the FM1 filter is ON.  This has been done for all the suspensions (BS, ITMX,ITMY, SRM, PRM, MC1, MC2,MC3, ITMX, ITMY).

I am also updating the first sensor filter banks for the BS, ITMX, ITMY, SRM, PRM,MC1,MC2,MC3, called "3:30", to match the Y and X ends.

8.  I can't find any documentation on how to get a momentary button press to toggle states.  I could stick a filter bank in and use the on/off feature of that part, but that feels like a silly hack.  I've decided for the moment to split the TM offset button into 2, one for ON, one for OFF.  I'll put in on the list of things to have added to the RCG code (either a method, or documentation if it already exists). 

EDIT:  TM offset still doesn't work.  Will worry about it next week.

9.  Fixed a connection in SPY/SPX models where  the side senor path that was missing a constant to a modulo block.

Quote:

I did some work on the ETMY real and Sim.

  1. Set SUS coil gains to have the same quadropole arrangement as the magnets do (-1, 1, 1, -1) so that POS = POS instead of pringle.
  2. Set the Sim Magnet polarities to match this. These are the ETMY_CI filter banks.
  3. Found that the Xycom cable from the ETMY slow controls was unplugged at the Xycom side. This was preventing enabling the ETMY coil driver and so there was no real damping of the suspension going on. I plugged it in and checked that the mirror could now be moved.
  4. The C1SUS_ETMY master screen's BLUE output filter area is now mis-labeled. If you trust the screen you would set it up to drive the suspension incorrectly. This MUST be fixed along with all of the other misleading features of the screen.
  5. ETMY SUSSIDE filter bank had a 2048 Hz sample rate and was making the damping not work correctly. Fixed to 16384 Hz.
  6. 12 Hz, 4th order Cheby low pass added and turned on for the local damping filtering. This is not optimum, but its just there to give us some filtering without introducing some instability via phase lag around 3 Hz.
  7. ETMY OL beam re-aligned on ISCT-EX.
  8. TM Offset buttons not working on the main overview screen.

It seems like there is still a problem with the input whitening filters. I believe the Xycom logic is set such that the analog whitening of the OSEM signals is turned ON only when the FM1 is turned OFF. Joe has got to fix this (and elog it) so that we can damp the suspension correctly. For now, the damping of the ETMY and the SETMY require different servo gains and signs, probably because of this.

 

ELOG V3.1.3-